OpenStudy (anonymous):

so if 7x = 2/7 what do i do next????

hero (hero):

Isolate x

hero (hero):

Most people would probably multiply both sides by 1/7: \[\left(\frac{1}{7}\right)7x = \frac{2}{7}\left(\frac{1}{7}\right)\]

hero (hero):

In this case you would end up with \[\frac{7x}{7} = \frac{(2)(1)}{(7)(7)}\]

hero (hero):

At this point you would cancel factors of one on the left side and multiply on the right side.